OpenStudy (momo_loves_cte):

4 1/5 - 2 2/5 = 2 -1/5

OpenStudy (momo_loves_cte):

thats not one of my answers.. what did i do wrong?

OpenStudy (anonymous):

\[=4\frac{ 1 }{ 5 }-2\frac{ 2 }{ 5 }\]

OpenStudy (anonymous):

\[=3\frac{ 6 }{ 5 }-2\frac{ 2 }{ 5 }\]

OpenStudy (anonymous):

6/5-2/5=?

OpenStudy (momo_loves_cte):

4/5

OpenStudy (anonymous):

yep

OpenStudy (momo_loves_cte):

So, its 1 4/5?

OpenStudy (anonymous):

Therefore the answer is\[=1\frac{ 4 }{ 5 }\]

OpenStudy (anonymous):

yup ^_^